# Reactive Knowledge Representation and Asynchronous Reasoning

### v1

@@ Abstract.
Background: Exact inference in complex probabilistic models often incurs prohibitive computational costs. This challenge is particularly acute for autonomous agents in dynamic environments that require frequent, real-time belief updates
Objectives: Existing methods are often inefficient for ongoing reasoning, as they re-evaluate the entire model upon any change, failing to exploit that real-world information streams have heterogeneous update rates. To address this, we approach the problem from a reactive, asynchronous, probabilistic reasoning perspective.
Methods: We first introduce Resin (Reactive Signal Inference), a probabilistic programming language that merges probabilistic logic with reactive programming. Furthermore, to provide efficient and exact semantics for Resin, we propose Reactive Circuits (RCs). Formulated as a meta-structure over Algebraic Circuits and asynchronous data streams, RCs are time-dynamic Directed Acyclic Graphs that autonomously adapt themselves based on the volatility of input signals.
Results: In high-fidelity drone swarm simulations, our approach achieves several orders of magnitude of speedup over frequency-agnostic inference. We demonstrate that RCs’ structural adaptations successfully capture environmental dynamics, significantly reducing latency and facilitating reactive real-time reasoning.
Conclusions: By partitioning computations based on the estimated Frequency of Change in the asynchronous inputs, large inference tasks can be decomposed into individually memoized sub-problems. This ensures that only the specific components of a model affected by new information are re-evaluated, drastically reducing redundant computation in streaming contexts.
copyright: ccdoi: 10.1613/jair.1.xxxxxjournalvolume: Xarticle: Ypublicationmonth: 1journalyear: 2026
@@ Acknowledgements.
Simon Kohaut gratefully acknowledges the financial support from the Honda Research Institute Europe (HRI-EU). The Eindhoven University of Technology authors received support from their Department of Mathematics and Computer Science and the Eindhoven Artificial Intelligence Systems Institute. This work was supported by the Konrad Zuse School of Excellence in Learning and Intelligent Systems (ELIZA). All emojis used in the figures were designed by OpenMoji under the CC BY-SA 4.0 license.

### v2

@@ Abstract.
Exact inference in complex probabilistic models often incurs prohibitive computational costs. This challenge is particularly acute for autonomous agents in dynamic environments that require frequent, real-time belief updates. Existing methods are often inefficient for ongoing reasoning, as they re-evaluate the entire model upon any change, failing to exploit that real-world information streams have heterogeneous update rates. To address this, we approach the problem from a reactive, asynchronous, probabilistic reasoning perspective. We first introduce Resin (Reactive Signal Inference), a probabilistic programming language that merges probabilistic logic with reactive programming. Furthermore, to provide efficient and exact semantics for Resin, we propose Reactive Circuits (RCs). Formulated as a meta-structure over Algebraic Circuits and asynchronous data streams, RCs are time-dynamic Directed Acyclic Graphs that autonomously adapt themselves based on the volatility of input signals. In high-fidelity drone swarm simulations, our approach achieves several orders of magnitude of speedup over frequency-agnostic inference. We demonstrate that RCs’ structural adaptations successfully capture environmental dynamics, significantly reducing latency and facilitating reactive real-time reasoning. By partitioning computations based on the estimated Frequency of Change in the asynchronous inputs, large inference tasks can be decomposed into individually memoized sub-problems. This ensures that only the specific components of a model affected by new information are re-evaluated, drastically reducing redundant computation in streaming contexts.
copyright: ccdoi: XXXXXXjournalvolume: Xarticle: 1publicationmonth: 1journalyear: 2026
@@ Acknowledgements.
Simon Kohaut gratefully acknowledges the financial support from the Honda Research Institute Europe (HRI-EU). The Eindhoven University of Technology authors received support from their Department of Mathematics and Computer Science and the Eindhoven Artificial Intelligence Systems Institute. This work was supported by the Konrad Zuse School of Excellence in Learning and Intelligent Systems (ELIZA). This work has benefited from the Cluster of Excellence ”Reasonable AI” funded by the German Research Foundation (DFG) under Germany’s Excellence Strategy— EXC-3057. All emojis used in the figures were designed by OpenMoji under the CC BY-SA 4.0 license.